OpenStudy (anonymous):

Find the area of the convex polygon in the plane with vertices at the points whose coordinates (-2,3), (1,10), (5,10), (8,7) and (4,0)... yes.. I have attempt it. I need help on it though! I graphed the shape and it looks like a house but diagonal on the first quadrant.

OpenStudy (anonymous):

Oh and it also passes the 2nd quadrant too.

OpenStudy (anonymous):

you break it into rectangles and triangles . the two triangles make a rectangle with an area \[\sqrt 42.5*\sqrt 86.5=\sqrt 3676.25\]the rectangle has an area 65. wheuh!!! That took forever!!!

OpenStudy (anonymous):

Thank you so much! You earned a medal for your hard work :) very much appreciate it!